painters A and B can paint a wall in 10 hours when working at the same time. Printer B works twice as fast as A.HOw long would it take each of them to paint if they worked alone?
Answer by rwm(914)

10/x+10/2x=1
multiply by 2x
20+10=2x
30=2x
15=x
Painter B can paint it in 15 hrs alone
painter A takes 30 hours alone
10/15+10/30=1
20/30+10/30=1
painter B does 2/3 while painter a does 1/3